x = 5410 , n=65, s=680 find 90% confidence interval for μ.


1
Expert's answer
2021-12-24T08:53:27-0500

The critical value for "\alpha = 0.1" and "df = n-1 = 64" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 1.669013."

The corresponding confidence interval is computed as shown below:


"CI=(x-t_c\times\dfrac{s}{\sqrt{n}}, x+t_c\times\dfrac{s}{\sqrt{n}})"

"=(5410-1.669013\times\dfrac{680}{\sqrt{65}},"

"5410+1.669013\times\dfrac{680}{\sqrt{65}})"

"=(5269.2294, 5550.7706)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "5269.2294 < \mu < 5550.7706," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(5269.2294, 5550.7706)."